OpenStudy (anonymous):

Write an expression for the perimeter and area of the rectangle with the given dimensions. I don't get how to do this. The width is 6 and the length is 4+2w

OpenStudy (anonymous):

ill answer this for 5 dollars :)

OpenStudy (anonymous):

I'm getting a cold and I can't figure this out. Can someone just give me an example problem? And I'll solve it?

OpenStudy (anonymous):

width = w = 6 length = L = 4+2w If we substitute the width into the length equation, we get L = 4+2*6 L= 16 Perimeter of a rectangle is 2w +2L = P 2*6 + 2*16 = P 12 + 32 = P 44 = P Area of a rectangle is L*w = Area 16*6 = Area 96 = Area

OpenStudy (anonymous):

Thanks senion. :)